OpenStudy (chelseatheruler):

-4y + 17 <= 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

\[-4y + 17 \le 29 \rightarrow -4y \le 29-17 \rightarrow -4y \le 12 \rightarrow y \le -3\] Thus the graph will pass through (0, -3)